OpenStudy (catelyncookies):

What is the equation of the graphed line? coordinates are : -6,-3 and 6,-7 then write in slope intercept form please show work thankyou :)

563blackghost (563blackghost):

You need to first find the slope \(\huge\bf{\frac{rise}{run} \rightarrow \frac{y_{2}-y_{1}}{x_{2}-x_{1}} \rightarrow \frac{-7+3}{6+6}=slope}\)